### v2

# Quantum Beatings in Optical Cavities

Abstract
Cavity polaritons, quasiparticles formed by coherent light-matter coupling, are at the heart of fundamental concepts of quantum optics. The quintessential signature of this coherent coupling is the Rabi oscillation, which results from the neglect of the counter-rotating-wave (CRW) effect in the weak-coupling regime. The goal of this letter is to predict resonant beatings that envelop the Rabi oscillation on the second or higher excitation manifold. These polariton beatings arise from the CRW term in the Dicke or Pauli-Fierz model and are directly correlated with the asymmetry in polariton eigenenergies. Our findings highlight the relevance of the CRW effect even in the weak-coupling regime, offer novel perspectives about coherent polariton dynamics, and shed new light on experiments of coupled quantum systems.
I Introduction
Cavity polaritons, quasiparticles arising from the coupling of molecules and photons in an optical cavity, have become a key concept in quantum optics.[1, 2] Along with superconducting circuits, trapped ions, and cold atoms in optical lattices, these hybrid quantum platforms have lent themselves to a diverse set of applications in quantum information, materials science, and condensed matter physics. Central to their theoretical descriptions are fundamental models describing the interaction between an ensemble of two-level systems (TLS) and the cavity photons, namely, the Tavis-Cummings (TC) model, the Dicke model (DM), and the Pauli-Fierz (PF) model. Common to these models is the prediction of Rabi oscillations, which are widely recognized as the quintessential signature of quantum coherence in the light-matter interaction.[3, 4] In this letter, we predict a more subtle quantum signature that goes beyond the Rabi oscillation in revealing the fine energy structure of light-matter hybrid states (i.e., polaritons) and in differentiating between various theoretical descriptions, even in the weak coupling regime.
II Model Hamiltonians
Using dimensionless units, the symmetrized TC model Hamiltonian[5] is given by
$H_{TC}=\omega_{m}J_{z}+\omega_{c}a^{{\dagger}}a+\frac{g}{\sqrt{N}}\left(J_{+}a+J_{-}a^{{\dagger}}\right)$ (1)
where $\omega_{m}$ and $\omega_{c}$ are the TLS and cavity frequencies, $N$ is the number of TLS, $g$ is the TLS-cavity coupling strength, $a$ and $a^{{\dagger}}$ are the creation and annihilation operators of the cavity mode, and $J_{\pm}$ are the collective raising and lowering operators for the symmetrized TLS. Notably, the interaction term in the TC Hamiltonian ensures the conservation of excitation number of the polariton system. Next, the Dicke Hamiltonian [6] is given by
$H_{DM}=\omega_{m}J_{z}+\omega_{c}a^{{\dagger}}a+\frac{g}{\sqrt{N}}J_{x}\left(a+a^{{\dagger}}\right)$ (2)
where $J_{x}=J_{+}+J_{-}$. Eq. (2) introduces the counter-rotating-wave (CRW) term, i.e. $H_{CRW}=g(J_{+}a^{\dagger}+J_{-}a)/\sqrt{N}$, which breaks the conservation of excitations. Finally, the PF Hamiltonian [7] is given by
$H_{PF}=H_{DM}+\frac{g^{2}}{\omega_{c}N}\left(J_{x}\right)^{2}$ (3)
where the additional term is the dipole self-energy (DSE) accounting for cavity-mediated interactions between the TLS. In the weak-coupling regime, the RWA is commonly used to simplify the analysis by neglecting the CRW term. The conditions for the RWA are $\omega_{c}-\omega_{m}=\Delta\omega\ll\omega_{c}+\omega_{m}$ (resonance or near-resonance) and $g\ll\omega_{m}$.
III Numerical Analysis
Numerical analysis of the cavity system involves the photon counts, spectral decomposition and energy schema. Simulations were carried out for $N$ from 1 to 15, but we focus primarily on the $N=2$ resonant case for the convenience of the subsequent perturbative analysis. The parameters are $\omega_{c}=\omega_{m}=1.0$, $0.01\leq g\leq 0.12$, and an initial state consisting of a TLS ensemble in the doubly-excited symmetric state and the cavity photon in the vacuum state, $\psi(0)=\ket{s_{2},0}$. To propagate the polariton system, we combine numerical integration of the Schrödinger equation and spectral decomposition [see Eq. (4)].
(a) TC Model
(b) Dicke Model
(c) PF Model
Figure 1: Average photon counts for all three models. Parameters: $g=0.07$, $\omega_{m}=\omega_{c}=1$, $N=2$.
III.1 Average Photon Count
The average photon count, i.e., the expectation value of photon number operator $a^{{\dagger}}a\right>$, is plotted for the TC, Dicke, and PF models in Fig. 1. These plots illustrate the presence of polariton beatings in the Dicke and PF models but not in the resonant TC model. Additionally, the beating dynamics in the Dicke and PF models are similar–the only noticeable difference between the two occurs between $t=1200$ and $t=1400$, where the oscillation amplitude diminishes. The similarity can be attributed to the DSE term in the PF model, which scales with $g^{2}/N$, deeming it an order of magnitude less significant than the CRW terms. Thus, we will focus on the Dicke model in the subsequent study of the beating dynamics. As shown in SM, a similar beating pattern can be observed on the higher excitation manifold. Further, the variance of the photon count also exhibits a complex beating pattern as shown in SM, suggesting quantum features in photon statistics.[8, 9]
Figure 2: The energy schema for the TC, Dicke, and PF models for $N=2$. The Hamiltonians are both scaled on-resonance with $g=0.07$. Note the agreement between the three schema for small $g$.
III.2 Spectral decomposition
The time-evolution of the polariton wave function can be evaluated via spectral decomposition,
$\psi(t)=\sum_{\lambda}\mid P_{\lambda}\rangle e^{-iE_{\lambda}t}\langle P_{\lambda}\mid\psi(0)\rangle=\sum_{\lambda}c_{\lambda}e^{-iE_{\lambda}t}\mid P_{\lambda}\rangle$ (4)
where $\psi(0)$ is the initial state and $c_{\lambda}=\langle P_{\lambda}\mid\psi(0)\rangle$ is the projection coefficient. In Eq. (4), $\lambda$ denotes a polariton eigen-state with the eigen-energy $E_{\lambda}$ and eigen-vector $P_{\lambda}$.
As suggested by Eq. (4), the beating can be caused by the eigenenergy $E_{\lambda}$ and eigenvector $P_{\lambda}$ as well as by the projection coefficient $c_{\lambda}$. Interestingly, reconstructing a wavefunction using TC eigenvalues with Dicke eigenvectors and Dicke projection coefficients yields virtually no beatings. Instead, using Dicke eigenvalues with TC eigenvectors and TC projection coefficients accurately reproduces the observed beating in the standard Dicke model (see Fig. 1 in Supplementary Material, i.e., SM). Thus, the quantum beatings in the Dicke model are caused by the energy shifts due to the CRW term and are well approximated by
$\psi(t)=\sum_{\lambda}c_{\lambda}e^{-iE_{\lambda}t}\mid P_{\lambda}\rangle\approx\sum_{\lambda}c^{TC}_{\lambda}e^{-iE^{DM}_{\lambda}t}\mid P^{TC}_{\lambda}\rangle$ (5)
where $c^{TC}$ and $P^{TC}$ are based on the TC model and $E^{DM}$ are the DM eigenenergies.
III.3 Energy Schema & Eigenvalue Asymmetry
The energy schema for the TC, Dicke, and PF models in Fig.  2 show the agreement between the three models for small $g$ and increasing energy shifts for large $g$. Note the symmetric structure in the TC model, the eigen-energy coalescence in the Dicke model, and the lower bound in the PF model.[6, 10, 11, 12, 13, 14, 15, 16] A less noticed feature is the eigenvalue asymmetry, defined as the difference between the middle polariton energy, $E_{0}$, and the average of the outer polariton energies, $(E_{+}+E_{-})/2$, i.e.,
$\alpha={E_{+}+E_{-}\over 2}-E_{0}.$ (6)
Evidently, the energy schema of the TC model is symmetric by construction, so $\alpha=0$ under the RWA in the TC model, but it grows with $g$ for the Dicke model, suggesting that the asymmetry $\alpha$ drives the quantum beating.
IV Photon count in the TC model
Within the second excitation manifold (SEM), we construct the symmetrized bright basis
$M_{2}^{s}=\{\ket{s_{0},2},\,\ket{s_{1},1},\,\ket{s_{2},0}\}$ (7)
where the first letter denotes the symmetric molecular state, with the subscript specifying the excitation level, and the second digit denotes the photon quantum number. Then, the SEM Hamiltonian becomes
$\begin{pmatrix}2\omega_{c}&\Omega_{N}&0\\ \Omega_{N}&\omega_{c}+\omega_{m}&\Omega_{N-1}\\ 0&\Omega_{N-1}&2\omega_{m}\end{pmatrix}$ (8)
where $\Omega_{M}=g\sqrt{2M/N}$. Solving this Hamiltonian in general is possible but tedious; we opt to consider the resonant case, $\omega_{m}=\omega_{c}$. [16] Then, we obtain three polariton solutions, $\{P_{+},P_{-},P_{0}\}$, with corresponding eigenvalues
$E_{0}=2\omega,\quad E_{\pm}=2\omega\pm\Omega$ (9)
where $\Omega=\sqrt{\Omega_{N}^{2}+\Omega_{N-1}^{2}}=\Omega_{2N-1}$ is the effective Rabi frequency of the SEM. Notably, the upper and lower polariton energies average to exactly the middle polariton energy such that $\alpha=0$. Next, we evaluate the time-dependence of the cavity photon expectation value and obtain
$n\right>(t)=-\frac{2(N-1)(1-4N+\cos(\Omega\,t))}{(2N-1)^{2}}\sin^{2}\left(\frac{\Omega\,t}{2}\right)$ (10)
which is derived in the SM. Eq. (10) predicts the Rabi oscillations in the TC model accurately.
(a) An outline of the perturbative approach. $M_{n}^{s}$ denotes the symmetric polariton states on the n-th excitation manifold.
(b) Beating period as a function of N, as predicted by Eq. 18. A local minimum is predicted between $N=9$ and $N=10$, and a divergence is predicted at N=5.
Figure 3:
V Perturbative analysis of beating
We proceed by deriving the beating dynamics and verifying the energy asymmetry hypothesis. To approximate an analytical form for the eigen-energy of the Dicke model, we perturb the TC eigen-energies by the CRW term.[16] First, we consider the coupling between the global ground state and the three polariton states $\{P_{+},P_{-},P_{0}\}$ on the SEM. The middle polariton $P_{0}$ is orthogonal to |$\ket{s_{1},1}$, so it is not affected. The upper and lower polaritons, $P_{\pm}$, are perturbed as
$\Delta E_{\pm}^{(0)}=\frac{\left|\braket{P_{\pm}|H_{CRW}|s_{0},0}\right|^{2}}{E_{\pm}}=\frac{g^{2}}{2(2\omega\pm\Omega)}$ (11)
which leads to accurate beating dynamics for N=2 (see SM).
Next, we consider the coupling to the fourth-excitation manifold (FEM) due to the CRW term $J_{+}a^{{\dagger}}$ To circumvent diagonalizing the TC Hamiltonian on the FEM, we adopt the product states instead of the hybrid states, which will shift the unperturbed energy $4\omega$ by the order of $\Omega$. Since $\Omega\propto g$, within the weak coupling regime, $\Omega/\omega$ is small, so we approximate all FEM eigenenergies by $4\omega$. Then, we write the perturbation on the SEM eigenenergies due to the coupling to the FEM as
$\Delta E_{i}^{(4)}=\sum_{n=1}^{3}\frac{|\braket{s_{n},4-n|J_{+}a^{{\dagger}}|P_{i}}|^{2}}{E_{i}-4\omega}$ (12)
where the resonance condition $E_{s_{1},3}=E_{s_{2},2}=E_{s_{3},1}=4\omega$ is used. As detailed in the SM, the FEM-perturbed energies are
$\Delta E_{0}^{(4)}=\frac{3g^{2}(3-2N)}{2\omega(2N-1)}$ (13)
$\Delta E_{\pm}^{(4)}=\frac{g^{2}(10+7N(2N-3))}{2N(2N-1)(-2\omega)}$ (14)
We then define $\Delta E_{i}=\Delta E_{i}^{(0)}+\Delta E_{i}^{(4)}$ such that the perturbed polariton eigenenergies in the Dicke model are
$E_{0}=2\omega+\Delta E_{0},\quad E_{\pm}=2\omega\pm\Omega+\Delta E_{\pm}$ (15)
Motivated by Eq. (5), we opt to keep the unperturbed TC eigenstates and projection coefficients, but adopt the perturbed DM eigenenergies. We make a further approximation by noticing that $|\Delta E_{-}^{(0)}-\Delta E_{+}^{(0)}|$ is generally small, so $\Delta E_{-}^{(0)}\approx\Delta E_{+}^{(0)}$. With these approximations, the average photon number is given by
$\left\langle n\right\rangle(t)=\frac{N-1}{2(2N-1)^{2}}\Bigg{[}\cos(2\,\Omega\,t)+8N-1$
 $\qquad-\,8N\,\cos(\alpha t)\,\cos(\Omega\,t)\Bigg{]}$ (16)
which reproduces the numerically exact results in the relevant parameter range. The above analysis allows us to identify the beating frequency as
$\alpha\approx{g^{2}\over 2\omega}{N-5\over N(2N-1)}$ (17)
which is a key result of this study.
Eq. (17) predicts the g and N dependence of the polariton beating. For example, the beating frequency scales quadratically with g, the Rabi frequency scales linearly, and thus their ratio $\alpha/\Omega$ is linear in g. As shown in Fig. 3 of the SM, the polariton beating becomes more accessible as g increases. Next, following Eq. (17), we obtain the beating period as
$T_{\text{beat}}={2\pi\over\alpha}\approx\frac{4\pi\omega}{g^{2}}\left|\frac{N(2N-1)}{N-5}\right|$ (18)
which predicts the maximum beating for $N=2$ and, interestingly, no beating for $N=5$. Numerical confirmation of this prediction is shown in Fig. 4, and the beating frequency is plotted as a function of $N$ in Fig. 3(b).
VI Experimental Relevance
Notably, modern experimental setups are capable of realizing the prediction of polariton beating. Consider a feasible physical system [17] with parameters $\frac{\omega_{c}}{2\pi}\approx 6$ GHz, $\frac{g}{2\pi}\approx 450$ MHz, which gives $\frac{g}{\omega_{c}}\approx 0.075$, precisely within the coupling range of our calculations. Further, we consider just two TLSs in the cavity; the Rabi oscillation period is then $T_{Rabi}\approx 1.56\text{ ns}$. Per Fig.  1, it takes roughly 25 coherent Rabi oscillations for a fully excited initial state to be maximally enveloped by the beating frequency. Experimentally realizing this would require a polariton coherence time approximately given by
${T_{\text{beat}}\over T_{\text{Rabi}}}\approx 25\implies T_{\text{coherent}}=T_{\text{beat}}\approx 39.3\text{ ns}$ (19)
The qubit decay rates and cavity decay rates of modern circuit QED systems allow for coherence times comfortably greater than 100 ns [18], demonstrating the relevance of our predictions. Experimental design of optical cavities, including those under vibrational strong coupling,[19, 20], can also approach this regime.
(a) $N=2$
(b) $N=3$
(c) $N=5$
Figure 4: Average photon count for the Dicke model: numerical results (blue) against the analytical approximations in Eq.  16 (red). Parameters: $g=0.07$, $\omega_{m}=\omega_{c}=1$. Note the lack of beating in plot (c).
VII Discussions
To explore the validity and implications of our prediction, we discuss the following issues: (1) off-resonant effects; (2) static disorder and dissipation; (3) quantum-classical correspondence.
  1. 1.
    The analysis presented in the letter is for the resonance case. Under the near-resonance condition, i.e. $\omega_{c}\neq\omega_{m}$ but $\omega_{c}-\omega_{m}\approx 0$, the TC model also causes noticeable beatings. Thus, the observed beating combines the contributions from the CRW term and detuning. As shown in the SM , we can differentiate their effects by their distinct sign dependence: The detuning contribution depends on the sign of detuning, whereas the CRW contribution is sign-independent. In fact, we can adjust the detuning to cancel the CRW contribution such that the beating vanishes.
  2. 2.
    The current calculation is based on the Hamiltonian description of isolated cavity polaritons and can be extended to incorporate various dissipative channels, including static disorder,[21, 22] stochastic noise (i.e. phonon couplings),[23] and photon loss.[24] For example,
    Further, a recent calculation of disordered polariton dynamics shows negligible effects on SEM polariton dynamics until the strength of static disorder reaches a critical value.[24] Therefore, it is reasonable to speculate that static disorder is perturbative below the critical value and various dissipative channels can be treated additively.
  3. 3.
    Finally, we discuss the implications of polariton beatings in the context of quantum-classical correspondence. As the textbook signature of light-matter coherence, Rabi oscillations or splittings have been studied within the semiclassical (e.g., Floquet) or mixed quantum-classical (MQC) framework. The polariton beating predicted in this paper is a subtle signature of light-matter coherence, and thus provides a stringent test of these approximations. The issue is worth further investigation given the extensive use of the MQC method in simulating complex molecular systems in optical cavities.[25, 26, 27] Along this line, an analogy can be established between polariton beatings analyzed here and anharmonic vibrations studied before.[28, 29, 30, 31] It stands to reason that Rabi oscillations are analogous to harmonic response whereas polariton beatings are analogous to anharmonic response, which requires more advanced treatments of quantum-classical correspondence.
VIII Conclusion
This letter predicts the emergence of quantum beatings in cavity polaritons. The characteristic oscillatory behavior arises from the CRW term in the Dicke or PF model and is thus absent in the resonant TC model. Our analysis demonstrates that the asymmetric shifts in the eigenenergies drive these beatings within the initial excitation manifold. The polariton beatings are fundamentally different from the Rabi oscillations, as they appear on longer time scales, vanish within the single-excitation manifold, and define a unique signature of the CRW effect. Experimental verification can be realized in optical lattices, superconducting circuits, and optical cavities including these under vibrational strong coupling.